OpenStudy (anonymous):

During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident? I'm just need help in finding the formula in the problem above so I can then solve it.

OpenStudy (anonymous):

is there 32 vehicles in total?

OpenStudy (anonymous):

Yes.

hero (hero):

Hint: If you let x = cars and y = trucks then 4x + 6y = 148 x + y =32

OpenStudy (anonymous):

Oh okay. I got the first part but couldn't think of the x+y=32. Oh okay Tyvm :)

OpenStudy (anonymous):

ok Fist put in random number, find the difference of should be and divide that number by 2 since the difference in truck wheel and car wheel is two and you need to add that number(the quotient )to car number and subtract the number from truck. If the number you got was smaller than the should be finde the difference, divide by two and add that to truck and subtract that number from car.

OpenStudy (anonymous):

In the problem by solving systems of equations by substituting I got y=10 and x= 22

hero (hero):

Yes, that is correct lol

hero (hero):

Good job